# Renormalon-like factorial enhancements to power expansion/OPE expansion in super-renormalizable QFTs

### v1

@@ 3.1.2 Mass insertion: $C_{n,1}$
To rule out the possibility that the renormalon above could cancel with the diagrams with mass-insertion, let’s consider $C_{n+1,1}$, namely coefficient functions with one mass insertions. There are three possibilities. Either the insertion is within the bubbles, or it is on the $p-k$ line, or it is on the external lines. Since the last type of contributions are trivial and factorizes out, here we only focuses on the first two-types of mass-insertions.
We first consider the mass insertion within the bubble chain. Notice that the mass-insertion diagram can be placed in $n$ locations in the bubble-chain with the same result, one can simply place it at the end of the bubble chain with a total number of $n$ bubbles. The diagram with one mass insertion within the bubble chain read
@@
$=\frac{m_{r}^{2}}{p^{4}}\left(\frac{\mu^{2}}{p^{2}}\right)^{\epsilon}\frac{(1+2\epsilon)f(\epsilon)}{2\pi\epsilon}\ .$ (124)
With the same $f(\epsilon)$ as before. Clearly, the partly-renormalized diagram reads
$R_{b}C_{n+1,1;\ a}\frac{m_{r}^{2}g^{n+1}}{(p^{2})^{n+1}}=$
@@
$+\frac{2nm_{r}^{2}g^{n+1}}{(4\pi)^{n}}\mu^{2\epsilon}\left(\frac{e^{\gamma_{E}}}{4\pi}\right)^{\epsilon}\frac{g(\epsilon)f(\epsilon)}{\epsilon^{n-1}}\int\frac{d^{d}k}{(2\pi)^{d}}\frac{1}{(k^{2})^{n+1}}\left(\frac{\mu^{2}}{k^{2}}\right)^{\epsilon}\bigg{(}f(\epsilon)\left(\frac{\mu^{2}}{k^{2}}\right)^{\epsilon}-1\bigg{)}^{n-1}\frac{1}{(p-k)^{2}}\ .$ (125)
Here $g(\epsilon)$ is again regular at $\epsilon=0$ and we have used the counter-term $a_{1}$ in Eq. (32). In fact, one has
$g(\epsilon)f(\epsilon)=2f(\epsilon)+\bigg{(}\frac{1}{\epsilon}-\frac{1}{2\epsilon}-\frac{1+\epsilon}{2\epsilon(1-\epsilon)}\bigg{)}\equiv 2f(\epsilon)-\frac{1}{(1-\epsilon)}\ .$ (126)
Also, we have separated the contributions into two terms. Also notice the overall factor of $n$. The second is actually a $n-1$ bubble chain. We can split the second term as
@@
Given the above, one can write
@@
Then one can calculate
$R_{b}C_{n+1,1;\ a}=$ $\frac{2n}{(4\pi)^{n+1}}\sum_{k=0}^{n}(-1)^{k}\binom{n}{k}\frac{G_{1}(n,\epsilon,(n+1-k)\epsilon)}{\epsilon^{n+2}(n+1-k)}$
$+\frac{2n}{(4\pi)^{n+1}}\sum_{k=0}^{n}(-1)^{k}\binom{n}{k}\frac{g(\epsilon)G_{1}(n,\epsilon,(n+1-k)\epsilon)}{\epsilon^{n+1}(n+1-k)}$
$+\frac{2n}{(4\pi)^{n+1}}\sum_{k=0}^{n-1}(-1)^{k}\binom{n-1}{k}\frac{g(\epsilon)G_{1}(n,\epsilon,(n-k)\epsilon)}{\epsilon^{n+1}(n-k)}\ .$ (129)
The $\frac{1}{\epsilon}$ term are generated only from the first and the last two terms, which are
@@
It is not hard to show that after adding the counter-term on the $p-k$ line, all the $\ln p$ in the $\frac{1}{\epsilon}$ cancel and lead to a $\frac{1}{\epsilon}$-independent operator counter-term. Given the above, using the same method as in the case of $C_{n,0}$ one obtains the leading factorial growth:
$C_{n,1;\ a}^{\rm asy1}=\frac{2n!}{(4\pi)^{n}}\frac{\mu^{2}}{p^{2}}-.....\ .$ (131)
@@
The mass-insertion on the $\frac{1}{(p-k)^{2}}$ line can also be calculated similarly. One has the partly renormalized bubble chain plus the counter-term on $p-k$
$R_{p}C_{n+1,1;\ b}\frac{m_{r}^{2}g^{n+1}}{(p^{2})^{n+1}}=$
@@
Notice that the last term is due to the subtraction of the $\Phi^{a}(i\partial)^{2}\Phi^{a}$ operator using the counter-term $a_{1}$ in Eq. (32). To show the cancellation of the $p^{2}$-dependent $\frac{1}{\epsilon}$ pole, one needs the relation
$-\frac{1}{n+1}\frac{1}{(p^{2})^{n+1}}\frac{d^{n+1}}{ds^{n+1}}\left(\frac{\mu^{2}}{p^{2}}\right)^{s}s(n+s)^{2}\bigg{|}_{s=0}$
$=-\left(\frac{\partial}{\partial p^{2}}+p^{2}\frac{\partial^{2}}{\partial(p^{2})^{2}}\right)\frac{1}{(p^{2})^{n}}\ln^{n}\frac{\mu^{2}}{p^{2}}\left(\frac{\mu^{2}}{p^{2}}\right)^{s}\bigg{|}_{s=0}\ .$ (133)
This implies that after subtracting all the IR sub-divergences, one only needs $p$-independent operator counter-terms. One can then proceed as the case of $C_{n,0}$ to obtain the factorial growths. The result reads
$C_{n,1;\ b}^{\rm asy1}=\frac{1}{(4\pi)^{n}}\frac{\mu^{2}}{p^{2}}n\times n!+....\ .$ (134)
@@
Clearly, not only it will not cancel the contribution in $C_{n,0}$, but it has an factor-$n$ enhancement. The enhancement is clearly due to the fact that the mass-insertion is roughly equivalent to two derivatives. One can proceeds further and show that the leading factorial growth to $C_{n,l}$, with all the mass-insertions on the $p-k$ line reads
$C_{n;l,0}^{\rm asy1}=-\frac{1}{(4\pi)^{n+1-l}}\frac{\mu^{2}}{p^{2}}\frac{(-1)^{l}\Gamma(n+1)^{2}}{\Gamma(l+1)^{2}\Gamma(n-l+1)}\ .$ (135)
@@
Here we introduced the new notation: $C_{n;l,k}$ denotes the contribution to $C_{n,l+k}$ where $l$ mass insertions are on the $p-k$ line, while $k$ mass insertions are within the bubble chain. Later we will see that the above summed over $k$ is almost the leading large-$n$ asymptotics of $C_{\rm I,NL}^{n}$, up to an over-all multiplication constant. For now, simply notice that the maximal value of $C_{n;l,0}^{\rm asy1}$ is achieved around $l=\sqrt{n}$, with is roughly $n!\frac{\sqrt{\frac{1}{n}}e^{2\sqrt{n}-\frac{1}{2}}}{2\pi}$.

### v2

@@ 3.1.2 Mass insertion: $C_{n,1}$
To rule out the possibility that the renormalon above could cancel with the diagrams with mass insertion, let’s consider $C_{n+1,1}$, namely coefficient functions with one mass insertions. There are three possibilities. Either the insertion is within the bubbles, or it is on the $p-k$ line, or it is on the external lines. Since the last type of contributions are trivial and has already been factorized out in our definition, here we only focus on the first two types of mass-insertions. In order to label the mass insertions, here we introduced the new notation: $C_{n;l,k}$ denotes the contribution to $C_{n,l+k}$ where $l$ mass insertions are on the $p-k$ line, while $k$ mass insertions are within the bubble chain. One clearly has $C_{n,k}=\sum_{l=0}^{k}C_{n;l,k-l}$.
Figure 9: The massless bubble chain diagram contributing to $C_{n+1;0,1}$, with one mass insertion within the bubbles for $n=5$ (upper) and the related IR subtractions (lower). Notice that the IR subtraction for the mass-inserted propagator is due to the $a_{1}$ in Eq. (32).
We first consider the mass insertion within the bubble chain, see Fig. 9 for a depiction. Notice that the mass-insertion diagram can be placed in $n$ locations in the bubble-chain with the same result, one can simply place it at the end of the bubble chain with a total number of $n$ bubbles. The bubble with one mass insertion reads
$C_{m}(k^{2})=-m_{r}^{2}\mu^{2\epsilon}\left(\frac{e^{\gamma_{E}}}{4\pi}\right)^{\epsilon}\int\frac{d^{D}l}{(2\pi)^{D}}\frac{1}{(k-l)^{2}}\frac{1}{(l^{2})^{2}}$
@@
$=\frac{m_{r}^{2}}{k^{4}}\left(\frac{\mu^{2}}{k^{2}}\right)^{\epsilon}\frac{(1+2\epsilon)f(\epsilon)}{2\pi\epsilon}\ ,$ (130)
with the same $f(\epsilon)$ as in Eq. (98). Here we need to perform the IR subtraction to $C_{m}$. Clearly, the $\frac{1}{(k-l)^{2}}$ propagator can still be re-normalized using the $\frac{1}{4\pi\epsilon}$ UV pole for $\frac{1}{N}\Phi^{2}$, while the $\frac{-m_{r}^{2}}{(l^{2})^{2}}$ propagator has to be renormalized by the UV pole for the $\frac{1}{N}\Phi(i\partial)^{2}\Phi$ operator insertion replacing the $\frac{1}{(l^{2})^{2}}$ line, using the Eq. (92) for the small-$l$ expansion rule of $\frac{1}{(k-l)^{2}}$ and Eq. (32) for the UV pole $a_{1}m_{r}^{2}$, which is exactly proportional to $m_{r}^{2}$. Therefore, the IR subtraction for the $C_{m}$ can be performed as
$\frac{m_{r}^{2}}{k^{4}}\left(\frac{\mu^{2}}{k^{2}}\right)^{\epsilon}\frac{(1+2\epsilon)f(\epsilon)}{2\pi\epsilon}+\frac{-m_{r}^{2}}{k^{4}}\times\frac{1}{4\pi\epsilon}+\frac{1}{k^{4}}\frac{1+\epsilon}{1-\epsilon}\times a_{1}m_{r}^{2}$
$=\frac{2m_{r}^{2}}{k^{4}}\frac{f(\epsilon)}{4\pi\epsilon}\left(f(\epsilon)\left(\frac{\mu^{2}}{k^{2}}\right)^{\epsilon}-1\right)+\frac{2m_{r}^{2}}{k^{4}}\frac{1}{4\pi}\left(\frac{\mu^{2}}{k^{2}}\right)^{\epsilon}2f(\epsilon)-\frac{2m_{r}^{2}}{k^{4}}\frac{1}{4\pi}g(\epsilon)\ ,$ (131)
where $g(\epsilon)$ is defined as
$g(\epsilon)=-\bigg{(}\frac{1}{\epsilon}-\frac{1}{2\epsilon}-\frac{1+\epsilon}{2\epsilon(1-\epsilon)}\bigg{)}\equiv\frac{1}{(1-\epsilon)}\ .$ (132)
It is again regular at $\epsilon=0$. Given the above, one can combine with the rest $n-1$ bubbles without the mass-insertion to obtain the bubble-chain-partly-renormalized diagram
$R_{b}\hat{C}_{n+1;0,1}\frac{m_{r}^{2}g^{n+1}}{(p^{2})^{n+1}}=$
@@
$+\frac{2nm_{r}^{2}g^{n+1}}{(4\pi)^{n}}\mu^{2\epsilon}\left(\frac{e^{\gamma_{E}}}{4\pi}\right)^{\epsilon}\frac{2f(\epsilon)}{\epsilon^{n-1}}\int\frac{d^{d}k}{(2\pi)^{d}}\frac{1}{(k^{2})^{n+1}}\left(\frac{\mu^{2}}{k^{2}}\right)^{\epsilon}\bigg{(}f(\epsilon)\left(\frac{\mu^{2}}{k^{2}}\right)^{\epsilon}-1\bigg{)}^{n-1}\frac{1}{(p-k)^{2}}$
$-\frac{2nm_{r}^{2}g^{n+1}}{(4\pi)^{n}}\mu^{2\epsilon}\left(\frac{e^{\gamma_{E}}}{4\pi}\right)^{\epsilon}\frac{g(\epsilon)}{\epsilon^{n-1}}\int\frac{d^{d}k}{(2\pi)^{d}}\frac{1}{(k^{2})^{n+1}}\bigg{(}f(\epsilon)\left(\frac{\mu^{2}}{k^{2}}\right)^{\epsilon}-1\bigg{)}^{n-1}\frac{1}{(p-k)^{2}}\ .$ (133)
We can further split the second term to obtain
@@
The bubble-chain-partly-renormalized $C_{n+1;0,1}$ therefore reads
$R_{b}\hat{C}_{n+1;0,1}=$ $\frac{2n}{(4\pi)^{n+1}}\sum_{k=0}^{n}(-1)^{k}\binom{n}{k}\frac{G_{1}(n,\epsilon,(n+1-k)\epsilon)}{\epsilon^{n+2}(n+1-k)}$
$+\frac{2n}{(4\pi)^{n+1}}\sum_{k=0}^{n}(-1)^{k}\binom{n}{k}\frac{2G_{1}(n,\epsilon,(n+1-k)\epsilon)}{\epsilon^{n+1}(n+1-k)}$
$+\frac{2n}{(4\pi)^{n+1}}\sum_{k=0}^{n-1}(-1)^{k}\binom{n-1}{k}\frac{(2-g(\epsilon))G_{1}(n,\epsilon,(n-k)\epsilon)}{\epsilon^{n+1}(n-k)}\ .$ (135)
Here the $G_{1}$ function is defined as
@@
The $\ln p$-dependent $\frac{1}{\epsilon}$ terms are generated only from the first and the last two terms in Eq. (3.1.2), which are
@@
It is not hard to show that after adding the counter-term $T_{p-k}R_{b}\hat{C}_{n+1;0,1}$ for the $p-k$ line, all the $\ln p$ in the $\frac{1}{\epsilon}$ cancel, leading to $\ln p$-independent poles for the operator . This time, the operator poles determined are the $g^{n+1}m_{r}^{2}$ terms for $O_{n+1}\rightarrow{\rm I}$ (or $g^{n}m_{r}^{2}$ term for $O_{n}\rightarrow{\rm I}$). The diagram for this operator pole can still be read from the upper of Fig. 8, in which $-\frac{Z_{n}g^{n+1}}{(4\pi)^{n+1}}$ should be replaced by the $m_{r}^{2}g^{n+1}$-order poles for the $O_{n+1}$. In fact, by staying at the same $n$ but considering all the $C_{n;0,k}$ with $10\leq k\leq n-$ 1, one determines all the operator poles for the $O_{n}\rightarrow{\rm I}$ mixing. One can proceed further using the same method as in the case of $C_{n,0}$ to obtain the fully renormalized $C_{n+1;0,1}$. Since this part of calculation is very similar to that of $C_{n,0}$, here we omit further details of it. The final result implies that there are also factorial growths in $C_{n;0,1}$. In fact, the leading non-alternating factorial growth reads:
$C_{n;0,1}\bigg{|}_{n\rightarrow\infty,+1}=\frac{2n!}{(4\pi)^{n}}\frac{\mu^{2}}{p^{2}}-.....\ .$ (138)
@@
Figure 10: The massless bubble chain diagram contributing to $C_{n+1;1,0}$, with one mass insertion one the $p-k$ line $n=5$ (upper) and the related IR subtractions (lower). Notice that the IR subtraction for the mass-inserted $p-k$ propagator is again due to the $a_{1}$ in Eq. (32). Notice that here although the IR counter-term $a_{1}$ do not correlate with IR counter-terms within the bubble chain, to perform the IR subtraction one must Taylor expand the IR subtracted bubble chain to the second order, see last term in Eq. (139).
The mass-insertion on the $\frac{1}{(p-k)^{2}}$ line can also be calculated similarly, see Fig. 10 for a depiction. This contribution is more illustrative than the mass-insertion within the bubble chain, since it requires the same operator poles as that of $O_{n}$ determined from the $C_{n,0}$ calculation, therefore it serves as a consistency check for the IR subtraction method. Here we can write directly the partly renormalized result (partly-renormalized-bubble-chain plus the IR counter-term on $p-k$)
$R_{p}\hat{C}_{n+1;1,0}\frac{m_{r}^{2}g^{n+1}}{(p^{2})^{n+1}}=$
@@
Notice that the last term is due to the subtraction of IR divergence in the small $p-k$ region using the UV pole $a_{1}$ in Eq. (32) for the $\frac{1}{N}\Phi(i\partial)^{2}\Phi$ operator. Clearly, in order to perform this subtraction, one must Taylor expand the IR subtracted bubble chain to the second order, explaining the second line of Eq. (139). To further simplify, one needs the function
$G_{2}(n,\epsilon,s)$ $=f(\epsilon)^{\frac{s}{\epsilon}-1}e^{\epsilon\gamma_{E}}\left(\frac{\mu^{2}}{p^{2}}\right)^{s}\frac{\epsilon\Gamma(-1-\epsilon)s\Gamma(1-n-s)\Gamma(1+n+s)}{\Gamma(-n-\epsilon-s)\Gamma(n+s-\epsilon)}$
$=\frac{(n+s)(n+s+\epsilon)}{1+\epsilon}G(n,\epsilon,s)\ .$ (140)
The pre-factor $\frac{(n+s)(n+s+\epsilon)}{1+\epsilon}$ suggests to write
$\frac{1}{(p^{2})^{n+1}}G_{2}(n,\epsilon,s)=\frac{1-\epsilon}{1+\epsilon}\left(\frac{\partial}{\partial p^{2}}+\frac{2p^{2}}{d}\frac{\partial^{2}}{\partial(p^{2})^{2}}\right)\frac{1}{(p^{2})^{n}}G(n,\epsilon,s)\ .$ (141)
In terms of the above, one has
$\frac{1}{(p^{2})^{n+1}}R_{p}\hat{C}_{n+1;1,0}=$ $-\frac{1-\epsilon}{1+\epsilon}\left(\frac{\partial}{\partial p^{2}}+\frac{2p^{2}}{d}\frac{\partial^{2}}{\partial(p^{2})^{2}}\right)\frac{1}{(p^{2})^{n}}R_{b}\hat{C}_{n,0}$
$-\left(\frac{\partial}{\partial p^{2}}+\frac{2p^{2}}{d}\frac{\partial^{2}}{\partial(p^{2})^{2}}\right)\frac{1}{(p^{2})^{n}}T_{p-k}R_{b}\hat{C}_{n,0}\ .$ (142)
Here $R_{b}\hat{C}_{n,0}$ is given in Eq. (3.1.1) and $T_{p-k}R_{b}\hat{C}_{n,0}$ is given in Eq. (3.1.1).
To proceed, one splits the second line of Eq. (3.1.2) to obtain
$\frac{1}{(p^{2})^{n+1}}R_{p}\hat{C}_{n+1;1,0}=$ $-\frac{1-\epsilon}{1+\epsilon}\left(\frac{\partial}{\partial p^{2}}+\frac{2p^{2}}{d}\frac{\partial^{2}}{\partial(p^{2})^{2}}\right)\frac{1}{(p^{2})^{n}}\left(R_{b}\hat{C}_{n,0}+T_{p-k}R_{b}\hat{C}_{n,0}\right)$
$-\frac{2\epsilon}{1+\epsilon}\left(\frac{\partial}{\partial p^{2}}+\frac{2p^{2}}{d}\frac{\partial^{2}}{\partial(p^{2})^{2}}\right)\frac{1}{(p^{2})^{n}}T_{p-k}R_{b}\hat{C}_{n,0}\ .$ (143)
From this equation, it is clear that the $\frac{1}{\epsilon}$ terms from the high-order derivatives again cancel. This is consistent with the fact that after subtracting all the IR sub-divergences, one only needs $p$-independent operator poles. In fact, the required operator poles here are still the $g^{n+1}$ terms in $Z_{n}$ for $O_{n}\rightarrow{\rm I}$, but attached to the $m_{r}^{2}$ power coefficient function:
$\frac{-Z_{n}}{(4\pi)^{n+1}}\times\frac{1}{(p^{2})^{n+1}}\frac{-n\Gamma(1-\epsilon)}{\Gamma(2+\epsilon)}\frac{\Gamma(n+\epsilon+1)}{\Gamma(n-\epsilon)}$
$=\frac{1-\epsilon}{1+\epsilon}\left(\frac{\partial}{\partial p^{2}}+\frac{2p^{2}}{d}\frac{\partial^{2}}{\partial(p^{2})^{2}}\right)\frac{1}{(p^{2})^{n}}\frac{Z_{n}}{(4\pi)^{n+1}}\frac{\Gamma(1-\epsilon)}{\Gamma(1+\epsilon)}\frac{\Gamma(n+\epsilon)}{\Gamma(n-\epsilon)}\ .$ (144)
See lower part of Fig. 8 for a depiction of this operator pole. Therefore, the contribution obtained by combining the operator poles and the first line of Eq. (143) in the $\epsilon\rightarrow 0$ limit, exactly leads to $C_{n,0}$ inside the derivatives
$-\frac{1-\epsilon}{1+\epsilon}\left(\frac{\partial}{\partial p^{2}}+\frac{2p^{2}}{d}\frac{\partial^{2}}{\partial(p^{2})^{2}}\right)\frac{1}{(p^{2})^{n}}\left(R_{b}\hat{C}_{n,0}+T_{p-k}R_{b}\hat{C}_{n,0}\right)$
$+\frac{1-\epsilon}{1+\epsilon}\left(\frac{\partial}{\partial p^{2}}+\frac{2p^{2}}{d}\frac{\partial^{2}}{\partial(p^{2})^{2}}\right)\frac{1}{(p^{2})^{n}}\frac{Z_{n}}{(4\pi)^{n+1}}\frac{\Gamma(1-\epsilon)}{\Gamma(1+\epsilon)}\frac{\Gamma(n+\epsilon)}{\Gamma(n-\epsilon)}$
$\rightarrow-\left(\frac{\partial}{\partial p^{2}}+p^{2}\frac{\partial^{2}}{\partial(p^{2})^{2}}\right)\frac{C_{n,0}}{(p^{2})^{n}}\ .$ (145)
Here we have used Eq. (114) for fully renormalized $C_{n,0}$. Therefore, using the $Z_{n}$ already given in Eq. (118), we have seen that it automatically renormalizes the coefficient function with one more mass insertion on the $p-k$ line. This is a crucial consistency check for our IR subtraction. On the other hand, the $\frac{2\epsilon}{1+\epsilon}$ term in the second line of Eq. (143) combines with the $\frac{1}{\epsilon}$ term in the $T_{p-k}R_{b}\hat{C}_{n,0}$ given in Eq. (3.1.1) to form a finite contribution.
Combining all above, the fully renormalized $C_{n+1;1,0}$ reads
$\frac{1}{(p^{2})^{n+1}}C_{n+1;1,0}$
$=-\left(\frac{\partial}{\partial p^{2}}+p^{2}\frac{\partial^{2}}{\partial(p^{2})^{2}}\right)\frac{C_{n,0}}{(p^{2})^{n}}+\frac{1}{(4\pi)^{n+1}}\left(\frac{\partial}{\partial p^{2}}+p^{2}\frac{\partial^{2}}{\partial(p^{2})^{2}}\right)\frac{2}{(p^{2})^{n}}\ln^{n}\frac{\mu^{2}}{p^{2}}\ .$ (146)
Clearly, the factorial growth is from the firs term. Using the previous result Eq. (128) for $C_{n,0}$, one therefore obtains the leading non-alternating asymtotics as
$C_{n;1,0}\bigg{|}_{n\rightarrow\infty,+1}=\frac{1}{(4\pi)^{n}}\frac{\mu^{2}}{p^{2}}n\times n!+....\ .$ (147)
@@
Clearly, not only it will not cancel the contribution in $C_{n,0}$, but it has a factor-$n$ enhancement in comparison to $C_{n,0}$. The enhancement is clearly due to the fact that the mass-insertion is roughly equivalent to two derivatives. One can proceed further without difficulty and show that the leading factorial growth of $C_{n;l,0}$, namely, all the $l$ mass-insertions are on the $p-k$ line, reads
$C_{n;l,0}\bigg{|}_{n\rightarrow\infty,+1}=-\frac{1}{(4\pi)^{n+1-l}}\frac{\mu^{2}}{p^{2}}\frac{(-1)^{l}\Gamma(n+1)^{2}}{\Gamma(l+1)^{2}\Gamma(n-l+1)}\ .$ (148)
@@
Later we will see that the above summed over $l$ is almost the leading large-$n$ asymptotics of $C_{\rm I,NL}^{n}$, up to an over-all multiplication constant $e^{-\frac{8\pi m_{r}^{2}}{g}}$ that is exponentially small in $g$. For now, simply notice that the maximal value of the asymptotics above is achieved around $l=\sqrt{n}$, with is roughly $n!\frac{\sqrt{\frac{1}{n}}e^{2\sqrt{n}-\frac{1}{2}}}{2\pi}$.